# Towards Hierarchical Multi-Agent Workflows for Zero-Shot Prompt Optimization

### v1

@@ Abstract
Large language models (LLMs) have shown great progress in responding to user questions, allowing for a multitude of diverse applications. Yet, the quality of LLM outputs heavily depends on the prompt design, where a good prompt might enable the LLM to answer a very challenging question correctly. Therefore, recent works developed many strategies for improving the prompt, including both manual crafting and in-domain optimization. However, their efficacy in unrestricted scenarios remains questionable, as the former depends on human design for specific questions and the latter usually generalizes poorly to unseen scenarios. To address these problems, we give LLMs the freedom to design the best prompts according to themselves. Specifically, we include a hierarchy of LLMs, first constructing a prompt with precise instructions and accurate wording in a hierarchical manner, and then using this prompt to generate the final answer to the user query. We term this pipeline Hierarchical Multi-Agent Workflow, or HMAW. In contrast with prior works, HMAW imposes no human restriction and requires no training, and is completely task-agnostic while capable of adjusting to the nuances of the underlying task. Through both quantitative and qualitative experiments across multiple benchmarks, we verify that despite its simplicity, the proposed approach can create detailed and suitable prompts, further boosting the performance of current LLMs.
@@ 1 Introduction
Large language models (LLMs) can perform a wide range of tasks such as creating detailed literary works and generating computer codes. Yet they often give unsatisfactory responses when the prompt is poorly designed. To design better prompts that make the LLM more effective, prompt optimization has been extensively explored in recent years. Specifically, existing works on prompt optimization can be categorized as follows. The first category involves the manual design of prompts, such as the chain of thought (CoT) [1] and its variants [2, 3, 4]. The second category focuses on optimizing prompts on a training set that contains input-response pairs to learn a golden prompt, as demonstrated by APE [5]. The third category also uses handcrafted prompts, but different from the first category, the handcrafted prompts include examples that can guide an LLM agent to further generate prompts for a second LLM that actually provide the answer, e.g., ExpertPrompting [6].
However, these methods generally have limited generalization performance across various tasks, as shown in Fig. 1 (a), (b), and (c). The first category [1] relies on handcrafted prompts. While they are very useful for tasks such as math, the same prompting method, if applied to other tasks, would be less effective. The second category, [5, 7], fine-tuned on a specific dataset, can obtain good performance for in-domain tasks. However, these learned prompts from one domain might fail for other tasks, making them not task-agnostic. The third category [5], which uses LLM to adjust the prompt, is still limited by the pre-defined examples in the handcrafted prompts. These examples only cover a finite number of scenarios and can potentially limit the quality of the modified prompt.
@@
On multiple benchmarks, we experimentally verify the effectiveness of the proposed hierarchical approach. When combined with Mixtral [8], HMAW achieves a significant average improvement of 30.7% across 5 datasets. The code for our work is available at here [^1] https://github.com/liuyvchi/HMAW
Figure 1: Examples comparing the generalization ability of existing methods and the proposed one. (a) COT [1] uses a handcrafted prompt, which might not be suitable for all tasks. (b) APE [5] fine-tunes the prompt on a specific dataset, and its generalization capability to other scenarios is questionable. (c) ExperPrompting [6] includes few-shot examples in the system prompt to help an LLM convert the user query to a format more suitable for LLM, but these examples might not be able to cover all scenarios. (d) Our method adopts a hierarchical design in reformatting the user query. Free from pre-defined few-shot examples, the interaction between the LLM hierarchy allows for more generalizable yet more adaptive tuning of the prompt.
@@ 2 Related Work
Manual prompt engineering integrates human-like problem-solving knowledge. For example, few-shot prompting [9] allows language models (LMs) to generate responses by providing them with explicit examples. The chain-of-thought (CoT) method [1], along with its variants such as zero-shot CoT [2], graph-of-thought (GoT) [3], and tree-of-thought (ToT) [4], intricately designs prompts to emulate various types of human-like reasoning processes. Some other works [10, 11, 12] manually craft prompts that encourage LLMs to engage in critical thinking and verification processes before delivering the final answer. Those manual prompt engineering techniques typically require carefully tailored prompts for various tasks, and for each task, the prompts used remain the same across different user queries. In comparison, we design a task-agnostic prompter that can create tailored prompts for every user query.
@@
      Chain of Thought (CoT)  [1]
      Zero-shot CoT  [2]
      Static ExpertPromting [6]
      Dynamic ExpertPromting [6]
      on Meta-Prompting [13]
      Rephrase and Response [14]
      Multi-Persona [15]
      APE [5]
      PromptAgent [16]
@@
Automated prompt engineering methods. Instead of using fixed hand crafted in-context prompts, automated prompt engineering methods are attracting increasing attention. This line of research aims to let LLMs be prompt generators which automatically generate specific prompts for each user query so that LLMs can respond better. Recent methods include automatic prompt engineer (APE) [5], expert prompting (EP) [6], synthetic prompting [17], skeleton-of-thought (SoT), [18], multi-personas [15], on-meta-prompting [13], meta-prompting [19], etc. While these methods can generate specific prompts according to different original prompts or instructions, they are generally not task-agnostic or zero-shot, where manually designed examples are required. In other words, they would not be very suitable for generalizing across ever changing landscape of LLMs. In comparison, our method is completely zero-shot, task-agnostic and prompt-specific.
Optimization based self-prompt refinement. Using optimization techniques to improve prompts is a recent focus. Early attempts including prompt tuning [20] and prefix tuning [21] tried to learn a trainable prefix or soft prompts for better prompting. Recent ones, e.g.,, directional stimulus prompting (DSP) [22], prompt-OIRL [23], use reinforcement learning to train an auxiliary model to generate query-specific prompts. OPRO [7] makes LLMs itself a optimizer which guides LLMs to produce outputs that align with a predefined objective. Those works typically require numerous iterations of training for each task and are thus not task-agnostic, and it is costly to adapt them to new tasks. Our work also optimizes user prompt and is thus closely related to this line of works. However, our method is learning-free and task-agnostic.
Zero-shot self-prompt refinement. Some manual prompt engineering methods, such as Zero-shot CoT [2], Static ExpertPrompting [6], and RaR [14], follow the zero-shot paradigm. Because no prompt examples or datasets are used, these methods are typically task-agnostic as well. Despite their simplicity and ease of adaptation to various tasks, their prompt optimization capability is limited by their shallow understanding of the user query. More importantly, they are not guaranteed to be query-specific (e.g., Static ExpertPrompting and RaR), which further limits their capability. In this work, we design a hierarchical multi-agent workflow to address these shortcomings.
@@ 3.1 Preliminaries
Some methods manually combine $q_{i}$ with some carefully designed fixed prefix or suffix to formulate $p_{i}^{*}$. For example, few-shot prompting [9] uses a few examples of query-answer pairs beside the query of interest $q_{i}$; the zero-shot chain-of-thought (CoT) [1] is formulated as ‘$q_{i}\oplus[\text{let us think step by step}]$’. Others use automatic methods, which are described in Section 2.
A recent line of research (e.g., RAR [14], ExpertPrompting [6], Multi-Persona [15]) focuses on leveraging the text generation capacity of LLMs to directly generate $p_{i}^{*}$ by providing some context $C\in\mathcal{T}$ of the optimization task in Eq. 1. Those methods are deployment-friendly because of characteristics like zero-shot, task-agnostic, and query-specific. In this paper, we not only keep those characteristics but also use a workflow to optimize prompts.
@@ 3.2 Hierarchical Multi-Agent Workflow
Intuition. Now, we consider prompt optimization as a complex and challenging task where prompts are hard to optimize by traditional optimization method. Our design is motivated by recent multi-agent workflow methods [24, 25] which have shown promise in solving complex tasks. We thus employ the idea of cooperative multiple agents in the prompt optimization task.
@@
Figure 2: Method Overview. We propose modeling the prompt optimization problem as a zero-shot output within a multi-agent workflow. The initial query, $q_{i}$, is first inputted into the first layer of our framework (the COE layer). Before being processed by the CEO LLM agent, $q_{i}$ is transformed into an LLM prompt $p_{i_{c}}$ by the prompter $f^{c}$, which also concatenates it with the context $C^{c}$ in the CEO layer. The output of the first layer, $q_{i}^{c}$, serves as the query from the CEO layer to the Manager layer. Similarly, the Manager Layer and the Worker Layer each include their own prompters, $f^{m}$ and $f^{w}$, respectively. Besides concatenating the content of this layer, the initial query $q_{i}$ is also concatenated to enhance stability. The input for the Worker LLM is our optimized prompt $P_{i}^{*}$, which directly triggers the LLM agent to generate the final response to the original query $q_{i}$.
@@ 4.1 Datasets and Evaluation Metrics
ATLAS [26] is a manually crafted benchmark for principled prompt evaluation. It contains a standard subset featuring questions across various domains, along with a challenging subset dedicated to reasoning and other complex tasks. It contains 520 questions, and each one has the corresponding output of various LLMs such as GPT-3.5 and Llama 2.
FED [27] is a comprehensive collection of annotated human-system and human-human conversations, featuring 18 fine-grained dialog qualities. It is comprised of 3,348 turn-level and 1,364 dialog-level data points, totaling 4,712. FED provides evaluation at both the turn and whole dialog levels. With moderate to strong correlation to human judgment across these levels, the dataset offers a reliable benchmark for analyzing interactive dialog systems.
GSM8K. The GSM8K dataset (Grade School Math 8K) [28] is a collection of 8,791 grade-school-level math word problems, with detailed step-by-step solutions. It includes 7,473 training problems and 1,319 test problems. Designed by OpenAI, GSM8K serves as a benchmark for evaluating mathematical reasoning in machine learning models. We report results on its test set.
CodeNet. Project CodeNet [29] is a large-scale dataset consisting of code samples across over 50 programming languages. We follow the practice in Self-Refine [30] and consider a random subset of 300 examples in Python. We compare different prompting optimization methods on the task of improving code readability.
@@
We use two metrics to evaluate the quality of generated responses. For the objective task GSM8K, we use the provided answer to compute the accuracy of the responses. For subjective tasks (i.e., Education, ATLAS, FED, CodeNet), we follow the practice in [30], which uses LLMs as the evaluator. Specifically, we compare the response pair generated from the optimized prompt and the response generated from the initial query by asking GPT-3.5 which response it prefers. GPT3.5 then assigns a score of 1 to the preferred response and a score of 0 to the other. The prompt used for the Evaluator agent is shown in the Appendix A.2. To reduce the impact of the primacy effect issue [31] and bias of LLMs in multiple choice questions, we follow the practice in PriDe [32] and permute option contents. Specifically, we run each test case twice with switched response orders and compute the average score. We then average these scores across the entire test set. The average score ranges from 0 to 1.
@@ 4.2 Compared Methods
Single Agent, Zero-shot. (1) No prompting. The initial query from the dataset is directly input into the LLMs without utilizing any additional prompting techniques. (2) Zero-shot CoT [1]. This method adds the fixed text ‘think step by step’ to the end of the initial query to formulate an optimized prompt for the LLM. (3) Rephrase and Response (RaR) [14]. This method allows LLMs to rephrase and expand questions posed by humans and provides responses in a single prompt.
Single Agent, Multi-shot (1) On Meta-prompting (OMP) [13]. It designs handcrafted meta-prompts to instruct LLMs to generate prompts for answering the initial human query. (2) Expert Prompting (EP) [6]: this method uses the LLM to generate a text description of the expert identity that is suitable to solve the given query, and then attaches the generated description to the initial query to create an optimized prompt.
Multi-agent prompting. Multi-Personas [15]. This method transforms a single LLM into a cognitive synergist by engaging in multi-turn self-collaboration with multiple personas.
Feedback-based Automatic Prompt Engineer (APE) [5]. This framework automatically generate prompt candidates and selete prompts based on LLM feedback on those candidates.
@@
Zero-CoT [1] 41.4 58.6 43.2 56.8 45.6 54.4 41.5 58.5 68.6 73.3 48.1 60.3 ($\uparrow$12.2)
RaR [14] 36.6 63.4 24.9 75.1 60.9 39.2 49.5 50.5 68.6 72.8 48.1 60.2 ($\uparrow$12.1)
On-MP [13] 53.6 46.4 31.8 68.2 50.7 49.3 47.5 52.5 68.6 61.3 50.4 55.5 ($\uparrow$5.1)
Static-EP [6] 44.0 56.0 39.2 60.8 42.4 57.6 52.5 47.5 68.6 73.5 49.3 59.1 ($\uparrow$9.8)
@@
Dynamic-EP [6] 45.5 54.5 68.6 71.3 44.8 55.3 46.9 53.1 68.6 66.9 54.9 60.2 ($\uparrow$5.4)
Multi-Persona [15] 73.2 26.8 50.9 49.1 75.2 24.8 57.5 42.5 68.6 71.5 65.1 42.9 ($\downarrow$22.2)

### v2

@@ Abstract
Large language models (LLMs) have shown great progress in responding to user questions, allowing for a multitude of diverse applications. Yet, the quality of LLM outputs heavily depends on the prompt design, where a good prompt might enable the LLM to answer a very challenging question correctly. Therefore, recent works developed many strategies for improving the prompt, including both manual crafting and in-domain optimization. However, their efficacy in unrestricted scenarios remains questionable, as the former depends on human design for specific questions and the latter usually generalizes poorly to unseen scenarios. To address these problems, we give LLMs the freedom to design the best prompts according to themselves. Specifically, we include a hierarchy of LLMs, first constructing a prompt with precise instructions and accurate wording in a hierarchical manner, and then using this prompt to generate the final answer to the user query. We term this pipeline Hierarchical Multi-Agent Workflow, or HMAW. In contrast with prior works, HMAW imposes no human restriction and requires no training, and is completely task-agnostic while capable of adjusting to the nuances of the underlying task. Through both quantitative and qualitative experiments across multiple benchmarks, we verify that despite its simplicity, the proposed approach can create detailed and suitable prompts, further boosting the performance of current LLMs. Project page: liuyvchi.github.io/HMAW_project/
@@ 1 Introduction
Large language models (LLMs) can perform a wide range of tasks such as creating detailed literary works and generating computer codes. Yet they often give unsatisfactory responses when the prompt is poorly designed. To design better prompts that make the LLM more effective, prompt optimization has been extensively explored in recent years. Specifically, existing works on prompt optimization can be categorized as follows. The first category involves the manual design of prompts, such as the chain of thought (CoT) (Wei et al., 2022) and its variants (Kojima et al., 2022; Besta et al., 2023; Yao et al., 2023). The second category focuses on optimizing prompts on a training set that contains input-response pairs to learn a golden prompt, as demonstrated by APE (Zhou et al., 2023). The third category also uses handcrafted prompts, but different from the first category, the handcrafted prompts include examples that can guide an LLM agent to further generate prompts for a second LLM that actually provide the answer, e.g., ExpertPrompting (Xu et al., 2023).
However, these methods generally have limited generalization performance across various tasks, as shown in Fig. 1 (a), (b), and (c). The first category (Wei et al., 2022) relies on handcrafted prompts. While they are very useful for tasks such as math, the same prompting method, if applied to other tasks, would be less effective. The second category, (Zhou et al., 2023; Yang et al., 2024a), fine-tuned on a specific dataset, can obtain good performance for in-domain tasks. However, these learned prompts from one domain might fail for other tasks, making them not task-agnostic. The third category (Zhou et al., 2023), which uses LLM to adjust the prompt, is still limited by the pre-defined examples in the handcrafted prompts. These examples only cover a finite number of scenarios and can potentially limit the quality of the modified prompt.
@@
On multiple benchmarks, we experimentally verify the effectiveness of the proposed hierarchical approach. When combined with Mixtral (Jiang et al., 2024), HMAW achieves a significant average improvement of 30.7% across 5 datasets.
Figure 1: Examples comparing the generalization ability of existing methods and the proposed one. (a) COT (Wei et al., 2022) uses a handcrafted prompt, which might not be suitable for all tasks. (b) APE (Zhou et al., 2023) fine-tunes the prompt on a specific dataset, and its generalization capability to other scenarios is questionable. (c) ExperPrompting (Xu et al., 2023) includes few-shot examples in the system prompt to help an LLM convert the user query to a format more suitable for LLM, but these examples might not be able to cover all scenarios. (d) Our method adopts a hierarchical design in reformatting the user query. Free from pre-defined few-shot examples, the interaction between the hierarchy allows for more generalizable yet more adaptive tuning of prompts.
@@ 2 Related Work
Manual prompt engineering integrates human-like problem-solving knowledge. For example, few-shot prompting (Brown et al., 2020) allows language models (LMs) to generate responses by providing them with explicit examples. The chain-of-thought (CoT) method (Wei et al., 2022), along with its variants such as zero-shot CoT (Kojima et al., 2022), graph-of-thought (GoT) (Besta et al., 2023), and tree-of-thought (ToT) (Yao et al., 2023), intricately designs prompts to emulate various types of human-like reasoning processes. Some other works (Welleck et al., 2023; Shinn et al., 2023; Gou et al., 2024) manually craft prompts that encourage LLMs to engage in critical thinking and verification processes before delivering the final answer. Those manual prompt engineering techniques typically require carefully tailored prompts for various tasks, and for each task, the prompts used remain the same across different user queries. In comparison, we design a task-agnostic prompter that can create tailored prompts for every user query.
@@
Chain of Thought (CoT)  Wei et al. (2022)
Zero-shot CoT  Kojima et al. (2022)
Static ExpertPromting (Xu et al., 2023)
Dynamic ExpertPromting (Xu et al., 2023)
on Meta-Prompting (de Wynter et al., 2023)
Rephrase and Response (Deng et al., 2023)
Multi-Persona (Wang et al., 2023b)
APE (Zhou et al., 2023)
PromptAgent (Wang et al., 2024)
@@
Automated prompt engineering methods. Instead of using fixed hand crafted in-context prompts, automated prompt engineering methods are attracting increasing attention. This line of research aims to let LLMs be prompt generators which automatically generate specific prompts for each user query so that LLMs can respond better. Recent methods include automatic prompt engineer (APE) (Zhou et al., 2023), expert prompting (EP) (Xu et al., 2023), synthetic prompting (Shao et al., 2023), skeleton-of-thought (SoT), (Ning et al., 2024), multi-personas (Wang et al., 2023b), on-meta-prompting de Wynter et al. (2023), meta-prompting (Suzgun & Kalai, 2024), etc. While these methods can generate specific prompts according to different original prompts or instructions, they are generally not task-agnostic or zero-shot, where manually designed examples are required. In other words, they would not be very suitable for generalizing across ever changing landscape of LLMs. In comparison, our method is completely zero-shot, task-agnostic and prompt-specific.
Optimization based self-prompt refinement. Using optimization techniques to improve prompts is a recent focus. Early attempts including prompt tuning Lester et al. (2021) and prefix tuning Li & Liang (2021) tried to learn a trainable prefix or soft prompts for better prompting. Recent ones, e.g.,, directional stimulus prompting (DSP) (Li et al., 2024), prompt-OIRL Sun et al. (2023), use reinforcement learning to train an auxiliary model to generate query-specific prompts. OPRO (Yang et al., 2024a) makes LLMs itself a optimizer which guides LLMs to produce outputs that align with a predefined objective. Those works typically require numerous iterations of training for each task and are thus not task-agnostic, and it is costly to adapt them to new tasks. Our work also optimizes user prompt and is thus closely related to this line of works. However, our method is learning-free and task-agnostic.
Zero-shot self-prompt refinement. Some manual prompt engineering methods, such as Zero-shot CoT (Kojima et al., 2022), Static ExpertPrompting (Xu et al., 2023), and RaR (Deng et al., 2023), follow the zero-shot paradigm. Because no prompt examples or datasets are used, these methods are typically task-agnostic as well. Despite their simplicity and ease of adaptation to various tasks, their prompt optimization capability is limited by their shallow understanding of the user query. More importantly, they are not guaranteed to be query-specific (e.g., Static ExpertPrompting and RaR), which further limits their capability. In this work, we design a hierarchical multi-agent workflow to address these shortcomings.
@@ 3.1 Preliminaries
Some methods manually combine $q_{i}$ with some carefully designed fixed prefix or suffix to formulate $p_{i}^{*}$. For example, few-shot prompting (Brown et al., 2020) uses a few examples of query-answer pairs beside the query of interest $q_{i}$; the zero-shot chain-of-thought (CoT) (Wei et al., 2022) is formulated as ‘$q_{i}\oplus[\text{let us think step by step}]$’. Others use automatic methods, which are described in Section 2.
A recent line of research (e.g., RAR (Deng et al., 2023), ExpertPrompting (Xu et al., 2023), Multi-Persona (Wang et al., 2023b)) focuses on leveraging the text generation capacity of LLMs to directly generate $p_{i}^{*}$ by providing some context $C\in\mathcal{T}$ of the optimization task in Eq. 1. Those methods are deployment-friendly because of characteristics like zero-shot, task-agnostic, and query-specific. In this paper, we not only keep those characteristics but also use a workflow to optimize prompts.
@@ 3.2 Hierarchical Multi-Agent Workflow
Intuition. Now, we consider prompt optimization as a complex and challenging task where prompts are hard to optimize by traditional optimization method. Our design is motivated by recent multi-agent workflow methods (Yang et al., 2024b; Wu et al., 2024) which have shown promise in solving complex tasks. We thus employ the idea of cooperative multiple agents in the prompt optimization task.
@@
Figure 2: Method Overview. We propose modeling the prompt optimization problem as a zero-shot output within a multi-agent workflow. The initial query, $q_{i}$, is first inputted into the first layer of our framework (the COE layer). Before being processed by the CEO LLM agent, $q_{i}$ is transformed into an LLM prompt $p_{i}^{c}$ by the prompter $f^{c}$, which also concatenates it with the context $C^{c}$ in the CEO layer. The output of the first layer, $q_{i}^{c}$, serves as the query from the CEO layer to the Manager layer. Similarly, the Manager Layer and the Worker Layer each include their own prompters, $f^{m}$ and $f^{w}$, respectively. Besides concatenating the content of this layer, the initial query $q_{i}$ is also concatenated to enhance stability. The input for the Worker LLM is our optimized prompt $P_{i}^{*}$, which directly triggers the LLM agent to generate the final response to the original query $q_{i}$.
@@ 4.1 Datasets and Evaluation Metrics
ATLAS (Bsharat et al., 2023) is a manually crafted benchmark for principled prompt evaluation. It contains a standard subset featuring questions across various domains, along with a challenging subset dedicated to reasoning and other complex tasks. It contains 520 questions, and each one has the corresponding output of various LLMs such as GPT-3.5 and Llama 2.
FED (Mehri & Eskenazi, 2020) is a comprehensive collection of annotated human-system and human-human conversations, featuring 18 fine-grained dialog qualities. It is comprised of 3,348 turn-level and 1,364 dialog-level data points, totaling 4,712. FED provides evaluation at both the turn and whole dialog levels. With moderate to strong correlation to human judgment across these levels, the dataset offers a reliable benchmark for analyzing interactive dialog systems.
GSM8K. The GSM8K dataset (Grade School Math 8K) (Cobbe et al., 2021) is a collection of 8,791 grade-school-level math word problems, with detailed step-by-step solutions. It includes 7,473 training problems and 1,319 test problems. Designed by OpenAI, GSM8K serves as a benchmark for evaluating mathematical reasoning in machine learning models. We report results on its test set.
CodeNet. Project CodeNet (Puri et al., 2021) is a large-scale dataset consisting of code samples across over 50 programming languages. We follow the practice in Self-Refine (Madaan et al., 2024) and consider a random subset of 300 examples in Python. We compare different prompting optimization methods on the task of improving code readability.
@@
We use two metrics to evaluate the quality of generated responses. For the objective task GSM8K, we use the provided answer to compute the accuracy of the responses. For subjective tasks (i.e., Education, ATLAS, FED, CodeNet), we follow the practice in (Madaan et al., 2024), which uses LLMs as the evaluator. Specifically, we compare the response pair generated from the optimized prompt and the response generated from the initial query by asking GPT-3.5 which response it prefers. GPT3.5 then assigns a score of 1 to the preferred response and a score of 0 to the other. The prompt used for the Evaluator agent is shown in the Appendix A.2. To reduce the impact of the primacy effect issue (Wang et al., 2023a) and bias of LLMs in multiple choice questions, we follow the practice in PriDe (Zheng et al., 2024) and permute option contents. Specifically, we run each test case twice with switched response orders and compute the average score. We then average these scores across the entire test set. The average score ranges from 0 to 1.
@@ 4.2 Compared Methods
Single Agent, Zero-shot. (1) No prompting. The initial query from the dataset is directly input into the LLMs without utilizing any additional prompting techniques. (2) Zero-shot CoT (Wei et al., 2022). This method adds the fixed text ‘think step by step’ to the end of the initial query to formulate an optimized prompt for the LLM. (3) Rephrase and Response (RaR) (Deng et al., 2023). This method allows LLMs to rephrase and expand questions posed by humans and provides responses in a single prompt.
Single Agent, Multi-shot (1) On Meta-prompting (OMP) (de Wynter et al., 2023). It designs handcrafted meta-prompts to instruct LLMs to generate prompts for answering the initial human query. (2) Expert Prompting (EP) (Xu et al., 2023): this method uses the LLM to generate a text description of the expert identity that is suitable to solve the given query, and then attaches the generated description to the initial query to create an optimized prompt.
Multi-agent prompting. Multi-Personas (Wang et al., 2023b). This method transforms a single LLM into a cognitive synergist by engaging in multi-turn self-collaboration with multiple personas.
Feedback-based Automatic Prompt Engineer (APE) (Zhou et al., 2023). This framework automatically generate prompt candidates and selete prompts based on LLM feedback on those candidates.
@@
Zero-CoT (Wei et al., 2022) 41.4 58.6 43.2 56.8 45.6 54.4 41.5 58.5 68.6 73.3 48.1 60.3 ($\uparrow$12.2)
RaR (Deng et al., 2023) 36.6 63.4 24.9 75.1 60.9 39.2 49.5 50.5 68.6 72.8 48.1 60.2 ($\uparrow$12.1)
On-MP (de Wynter et al., 2023) 53.6 46.4 31.8 68.2 50.7 49.3 47.5 52.5 68.6 61.3 50.4 55.5 ($\uparrow$5.1)
Static-EP (Xu et al., 2023) 44.0 56.0 39.2 60.8 42.4 57.6 52.5 47.5 68.6 73.5 49.3 59.1 ($\uparrow$9.8)
@@
Dynamic-EP (Xu et al., 2023) 45.5 54.5 68.6 71.3 44.8 55.3 46.9 53.1 68.6 66.9 54.9 60.2 ($\uparrow$5.4)
Multi-Persona (Wang et al., 2023b) 73.2 26.8 50.9 49.1 75.2 24.8 57.5 42.5 68.6 71.5 65.1 42.9 ($\downarrow$22.2)